OpenStudy (anonymous):

NEED HELP PLEASE ON MATH The populations of six countries are listed in this chart. Arrange the population of the six countries from least to greatest. Explain how you solved this problem.

OpenStudy (anonymous):

Country Population (number of people) Japan = 12.718 × 10^7 Brazil= 2.0 × 10^8 Germany= 80.619 × 10^6 United Kingdom= 63,708,000 Nigeria= 0.00173615 × 10^11 Bangladesh= 15251 × 10^4

OpenStudy (anonymous):

can anyone help??